# Creating General User Models from Computer Use

### v2

@@ 4.3.1. Discovering Suggestions
Proposition: [Anon] is likely going to friend’s wedding in Chicago.
@@
Proposition: [Anon] doesn’t own any suitable formal wear.
@@
Proposition: [Anon] needs to budget as a Ph.D. student.
@@ 4.3.2. Determining if suggestions are worth showing
When necessary, suggestions generated by Gumbo can be delegated to a set of tools or agents. We use an additional zero-shot prompt to determine if a suggestion requires a tool (see Appendix B.2.2). We implement a handful of external tools which can be toggled by the user (and are by default disabled) to preserve privacy. Gumbo delegates calls to Gemini 2.0 Flash only if search with Google or code execution is required. In addition, Gumbo implements file system search and retrieval through the local MacOS spotlight search, and computer use via the Operator API. [^2] We leave Computer Use disabled in Gumbo’s evaluation, both to reduce external dependencies and since we found computer use too slow / buggy. We expect advancements in computer use to greatly benefit from implementing GUMs
@@ 5.2. Auditing Observations
[Anon] is careful about sharing access credentials
[Anon] has discussed privacy concerns with students in the past
[Anon] only discusses financial matters with his spouse
@@ 5.3. Constructing Propositions
Once the proposition and rationale are generated, we use GUM’s underlying LLM to produce a confidence score in the proposition. Confidence scores support future revisions of uncertain propositions and help applications make decisions under uncertainty. While we could, in theory, use the LLM’s own log probabilities by looking directly at the logit scores on the completion , these predictions are often overly confident. Like Tian et al. (2023), we observe that simply prompting the model to generate confidence scores yields more calibrated outputs instead of looking at underlying logits. [^3] Our setting is slightly different: we’re looking at confidence in a user’s beliefs, instead of confidence in the model’s own response—but Tian et al. (2023) generalizes. We prompt the model to generate a confidence score on a 1-10 scale, and then normalize between 0-1. Our full prompt is in Appendix B.1.1.
@@
Some propositions remain more relevant over time compared to others. Unfortunately for the first author, a proposition like ‘‘User is eating ice cream.’’ decays faster than ‘‘User is Ph.D. student.’’ (Such is the nature of life. But, a proposition such as ‘‘User enjoys eating ice cream more than their Ph.D.’’ [^4] Yes, a real proposition generated by our system.—that’s timeless.) To account for this, we take inspiration from cognitive architectures (Anderson, 1993; Park et al., 2023) and introduce a decay factor. Instead of applying a global decay to all propositions universally, we use the underlying LLM to generate a proposition-specific decay. We include all information about the proposition in-context (reasoning and confidence) and generate a decay score from 1-10, e.g. $\mathrm{decay}\sim P_{LM}(\cdot\mid\mathrm{confidence},\mathrm{proposition},\mathrm{reasoning},\mathrm{obs)}$.
@@ 5.4. Retrieving Context-Sensitive Propositions
The process so far produces propositions and confidence scores (from Propose, §5.3), but we need to compare these with what GUM already knows. To enable comparison and retrieval over GUMs, we implement a Retrieve module. Speed is critical: each observation generates multiple propositions that require hundreds of queries. We use a two-stage retrieve and rerank approach (Nogueira and Cho, 2019). First, we retrieve documents using BM25, which considers term frequency, inverse document frequency, and document length. [^5] While we use lexical similarity, one can easily replace our approach with neural embeddings Then, we use an LLM-based re-ranker to improve precision. Our setup optimizes for speed but can accommodate other retrieval/reranking methods.
@@
BM25 computes relevance scores, but recent propositions should be prioritized. Each proposition $d_{i}$ includes a decay score $\alpha_{i}$5.3). We define the decay factor as $\gamma_{i}=\exp\bigl{(}-\alpha_{i}\cdot k\cdot\text{age}(d_{i})\bigr{)}$, where $k=2$ and $\text{age}(d_{i})$ is a continuous value measured in days. The adjusted relevance score is:
@@ 6.1. GUM Ablations
  • Ablation: No {Retrieve, Revise}. In this ablation, we test a version of GUM that only creates propositions with no confidence. The GUM is unable to retrieve past context—generations are constructed given only the current email.
@@ 6.3. Evaluation Methods
Because GUM generates confidence scores on a $1-10$ scale, we normalize these scores to the [0,1] range before computing the Brier score. Finally, we conducted t-tests for Brier across GUM ablations.
@@ 6.4. Results
To differentiate between ablations and to identify relative performance of the full GUM, we asked participants to rank ablations. In the ranked setting, the difference between the full GUM the ablations are clear (Fig 5). Win rates for the full model are higher than the ablations ($\mu$, 95% conf; Full = $0.618\pm 0.07$ , No {Retrieve} $=0.463\pm 0.07$, No {Retrieve, Revise} $=0.420\pm 0.07$). This difference is also validated by a bootstrapped significance test following the Holm-Beniforri correction ($p<0.05$). Between the No {Retrieve, Revise} and No {Revise} ablations, however, we see no significant difference. We suspect that relevant email is often spaced apart, and that revision without retrieval (e.g. the sliding window approach) fails to capture long-range dependencies between emails. Without this, Revision falls apart entirely.
@@
On average, participants rank GUM propositions as correct $76.15\pm 5.85$ of the time. Participants were also fairly positive on the general accuracy of propositions.
@@ 8.1. Procedure
8.1.1. Participants and Infrastructure
We use the same infrastructure in our technical evaluation, except that we deploy a vision model (Qwen 2.5 VL 72B (Bai et al., 2025)) for the Screen Observer, transcribing observations from the screen. Participants were recruited using a mix of snowball sampling and word-of-mouth. Given the level of trust required to use a system like Gumbo, participants were either acquaintances of the authors or acquaintances of acquaintances. No participants had any knowledge of the system before the study. Participants were located in the United States and were over 18 years of age. 2 participant identified as White, 2 as Asian, and 1 as African American, and all had bachelor’s degrees. Participation required about an hour for synchronous onboarding and offboarding; compensation was \$75 USD.
@@ 8.3. Results
Mirroring results from our email evaluation, participants found sampled propositions to be accurate ($0.79\pm 0.07$) and calibrated (Brier $=0.28\pm 0.04$), with error coming from underconfident predictions. Propositions that reflected factual statements about the user were often annotated or discussed with very little hesitation.
@@ 8.4. Errors and Boundaries
(Is) the screenome limited? Does it generalize?
@@ 9.4. Limitations and Future Work
Better models will extend beyond vision and language
The current GUM implementation relies on vision and language models, constraining the input space. Large models that ingest a wider range of modalities  (Bommasani et al., 2021) and diverse tool uses could enhance GUM’s capabilities. Already, advanced speech models can discern a user’s tone, while healthcare models interpret various sensor signals—all of which in principle could integrate into the GUM.

### v3

@@ 4.3.1. Discovering Suggestions
Proposition: Omar is likely going to friend’s wedding in Chicago.
@@
Proposition: Omar doesn’t own any suitable formal wear.
@@
Proposition: Omar needs to budget as a Ph.D. student.
@@ 4.3.2. Determining if suggestions are worth showing
When necessary, suggestions generated by Gumbo can be delegated to a set of tools or agents. We use an additional zero-shot prompt to determine if a suggestion requires a tool (see Appendix B.2.2). We implement a handful of external tools which can be toggled by the user (and are by default disabled) to preserve privacy. Gumbo delegates calls to Gemini 2.0 Flash only if search with Google or code execution is required. In addition, Gumbo implements file system search and retrieval through the local MacOS spotlight search, and computer use via the Operator API. [^2] We leave Computer Use disabled in Gumbo’s evaluation, both to reduce external dependencies and since we found computer use too slow / buggy. We expect advancements in computer use to greatly benefit from implementing GUMs.
@@ 5.2. Auditing Observations
Michael is careful about sharing access credentials
Michael has discussed privacy concerns with students in the past
Michael only discusses financial matters with his spouse
@@ 5.3. Constructing Propositions
Once the proposition and rationale are generated, we use GUM’s underlying LLM to produce a confidence score in the proposition. Confidence scores support future revisions of uncertain propositions and help applications make decisions under uncertainty. While we could, in theory, use the LLM’s own log probabilities by looking directly at the logit scores on the completion, out of the box model predictions are often overly confident. Like Tian et al. (2023), we observe that simply prompting the model to generate confidence scores yields more calibrated outputs instead of looking at underlying logits. [^3] Our setting is slightly different: we’re looking at confidence in a user’s beliefs, instead of confidence in the model’s own response—but Tian et al. (2023) generalizes. We prompt the model to generate a confidence score on a 1-10 scale, then normalize between 0-1. Our full prompt is in Appendix B.1.1.
@@
Some propositions remain more relevant over time compared to others. Unfortunately for the first author, a proposition like ‘‘User is eating ice cream.’’ decays faster than ‘‘User is Ph.D. student.’’ (Such is the nature of life. But, a proposition such as ‘‘User enjoys eating ice cream more than their Ph.D.’’ [^4] Yes, a real proposition generated by our system.—that’s timeless.) To account for this, we take inspiration from cognitive architectures (Anderson, 1993; Park et al., 2023) and introduce a decay score. Instead of applying a global decay to all propositions universally, we use the underlying LLM to generate a proposition-specific decay. We include all information about the proposition in-context (reasoning and confidence) and generate a decay score from 1-10, e.g. $\mathrm{decay}\sim P_{LM}(\cdot\mid\mathrm{confidence},\mathrm{proposition},\mathrm{reasoning},\mathrm{obs)}$.
@@ 5.4. Retrieving Context-Sensitive Propositions
The process so far produces propositions and confidence scores (from Propose, §5.3), but we need to compare these with what GUM already knows. To enable comparison and retrieval over GUMs, we implement a Retrieve module. Speed is critical: each observation generates multiple propositions that require hundreds of queries. We use a two-stage retrieve and rerank approach (Nogueira and Cho, 2019). First, we retrieve propositions and underlying observations (our “document”) using BM25, which considers term frequency, inverse document frequency, and document length. [^5] While we use lexical similarity, one can easily replace our approach with neural embeddings Then, we use an LLM-based re-ranker to improve precision. Our setup optimizes for speed but can accommodate other retrieval/reranking methods.
@@
BM25 computes relevance scores, but recent propositions should be prioritized. Each proposition $d_{i}$ includes a decay score $\alpha_{i}$5.3). We define the decay score as $\gamma_{i}=\exp\bigl(-\alpha_{i}\cdot k\cdot\text{age}(d_{i})\bigr)$, where $k=2$ and $\text{age}(d_{i})$ is a continuous value measured in days. The adjusted relevance score is:
@@ 6.1. GUM Ablations
  • Ablation: No {Retrieve, Revise}. In this ablation, we test a version of GUM that only creates propositions with no confidence. The GUM is unable to retrieve past context—propositions are constructed given only the current email.
@@ 6.3. Evaluation Methods
Because GUM generates confidence scores on a 1101-10 scale, we normalize these scores to the [0,1] range before computing the Brier score. Finally, we conducted t-tests for Brier across GUM ablations.
@@ 6.4. Results
To differentiate between ablations and to identify relative performance of the full GUM, we asked participants to rank output propositions from each ablation. In the ranked setting, the difference between the full GUM and the ablations are clear (Fig 5). Win rates for the full model are higher than the ablations ($\mu$, 95% conf; Full = $0.618\pm 0.07$ , No {Retrieve} $0.463\pm 0.07$, No {Retrieve, Revise} $0.420\pm 0.07$). This difference is also validated by a bootstrapped significance test following the Holm-Beniforri correction ($p<0.05$). Between the No {Retrieve, Revise} and No {Revise} ablations, however, we see no significant difference. We suspect that relevant email is often spaced apart, and that revision without retrieval (e.g. the sliding window approach) fails to capture long-range dependencies between emails. Without this, Revision falls apart entirely.
@@
Highly confident (confidence >0.8>0.8) propositions were labeled as correct $88.2\%$ of the time by participants. When confidence =1.0=1.0, propositions are completely correct ($100\%$ accuracy). Even when we include all propositions, regardless of confidence score, participants rank GUM propositions as correct $76.15\pm 5.85$ of the time—lower, but still fairly accurate. Participants were also fairly positive on the general accuracy of propositions.
@@
ID Context Role / Program Representative Computer Activities
P1 Work & Personal HCI graduate student Reading news; writing a paper; communicating with friends and family
P2 Work & Personal HCI graduate student Brainstorming; personal communication
P3 Work Civil engineering graduate student Brainstorming; qualifying-exam preparation; reading related work
P4 Personal Data scientist Watching NCAA games; wedding planning; looking for a new job
P5 Work & Personal NLP graduate student Data processing; gaming (Twitch streams)
Table 2. Overview of Participant Roles and Computer Activities
@@ 8.1. Procedure
8.1.1. Recruitment and Infrastructure
We use the same infrastructure in our technical evaluation, except that we deploy a vision model (Qwen 2.5 VL 72B (Bai et al., 2025)) for the Screen Observer, transcribing observations from the screen. Participants were recruited using a mix of snowball sampling and word-of-mouth. Given the level of trust required to use a system like Gumbo, participants were either acquaintances of the authors or acquaintances of acquaintances. Participation required about an hour for synchronous onboarding and offboarding; compensation was \$75 USD.
8.1.2. Participant Details
No participants had any knowledge of the system before the study. Participants were located in the United States and were over 18 years of age. 2 participant identified as White, 2 as Asian, and 1 as African American, and all had bachelor’s degrees. All participants reported what they primarily used their computers for (work or personal) along with a high level list of activities (see Table  2).
@@ 8.3. Results
Mirroring results from our email evaluation, participants found sampled propositions to be accurate ($0.79\pm 0.07$) and calibrated (Brier $0.28\pm 0.04$), with error coming from underconfident predictions. Propositions that reflected factual statements about the user were often annotated or discussed with very little hesitation.
@@ 8.4. Errors and Boundaries
(Is) screen activity limited? Does it generalize?
@@ 9.4. Limitations and Future Work
Better models will support longer contexts, be smaller, and extend beyond vision and language
Many components of the GUM pipeline work to mitigate the limitations of current-day LLMs, as simply putting everything in context is insufficent. Models forget parts of the context (Liu et al., 2023) or yield generic propositions. Our retrieval step, for example, is required as models cannot reliably process extremely long context lengths. Explicit confidence verbalization is necessary as model logits are uncalibrated. However, we expect future work on modeling for GUMs (e.g. end-to-end learning parts of our pipeline) to subsume specific steps. We also expect future models to grow smaller in size and run entirely on-device, further mitigating privacy risks while reducing storage impact. Currently, the GUM saves about  1GB of screenshots per day (just 5 MB of transcribed observations), which will accrue if deployed for longer. Another avenue for future work involves collecting minimal observational data for effective user models. Finally, the current GUM implementation relies on vision and language models, constraining the input space. Large models that ingest a wider range of modalities  (Bommasani et al., 2021) along with diverse tool use could enhance GUM’s capabilities. Already, advanced speech models can discern a user’s tone, while healthcare models interpret various sensor signals—all of which in principle could integrate into the GUM.
@@ 10. Conclusion
Acknowledgements.
We thank Michael Y. Li, Jensen Gao, Suvir Mirchandani, Lindsay Popowski, Ryan Louie, Beleicia Bullock, Will Held, Dora Zhao, Tiziano Piccardi, Michelle Lam, and Carolyn Zou for helpful discussions and feedback. We also thank members of the Stanford HCI and NLP groups for feedback in general and for testing both GUM and Gumbo. We don’t thank Omar’s ice cream making hobby for filling this paper with examples that might make the reader hungry. Omar Shaikh is supported by the HAI-HPI program and Joon Park is supported by the Microsoft Research fellowship. We also thank SCBx, AXA, TRI, American Express, Banco Itaú, Ford, Hanwha, and Google, as well as ONR grant N000142412532 for helping fund our work.